OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = –3x + 12 and passes through the point (−1, 6). A y = 1 over 3x + 7 B y = –3x + 3 C y = 1 over 3x + 3 D y = –3x + 7 My answer is B am I correct?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

@dread907 thanks, could you help me with another one please

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = 3 over 2x + 4 and passes through the point (−6, 3). A y = 3 over 2x − 1 B y = 3 over 2x + 3 C y = −2 over 3x − 1 D y = −2 over 3x + 3

OpenStudy (anonymous):

What do you think it is first, and why? Walk yourself through it and do out the equation.

OpenStudy (anonymous):

D @dread907 is that correct

OpenStudy (anonymous):

Yes, that is correct good job!